One of Shakespeare's sonnets has a verb in 13 of its 19 lines, an adjective in 12 lines, and both in 9 lines.
How many lines have a verb but no adjective?

How many lines have an adjective but no verb?

How many have neither an adjective nor a verb?

Answer by ikleyn(52781)

.
One of Shakespeare's sonnets has a verb in 13 of its 19 lines, an adjective in 12 lines, and both in 9 lines.

How many lines have a verb but no adjective?              13 - 9 = 4
 
How many lines have an adjective but no verb?             12 - 9 = 3

    How many lines have EITHER an adjective OR a verb ?   13 + 12 - 9 = 25 - 9 = 16
 
How many have neither an adjective nor a verb?            19 - 16 = 3


To learn more and to learn more details, see the leson
    - Counting elements in sub-sets of a given finite set
in this site.